# Data Distributional Properties As Inductive Bias for Systematic Generalization

### v2

@@ Abstract
Deep neural networks (DNNs) struggle at systematic generalization (SG). Several studies have evaluated the possibility to promote SG through the proposal of novel architectures, loss functions or training methodologies. Few studies, however, have focused on the role of training data properties in promoting SG. In this work, we investigate the impact of certain data distributional properties, as inductive biases for the SG ability of a multi-modal language model. To this end, we study three different properties. First, data diversity, instantiated as an increase in the possible values a latent property in the training distribution may take. Second, burstiness, where we probabilistically restrict the number of possible values of latent factors on particular inputs during training. Third, latent intervention, where a particular latent factor is altered randomly during training. We find that all three factors significantly enhance SG, with diversity contributing an 89% absolute increase in accuracy in the most affected property. Through a series of experiments, we test various hypotheses to understand why these properties promote SG. Finally, we find that Normalized Mutual Information (NMI) between latent attributes in the training distribution is strongly predictive of out-of-distribution generalization. We find that a mechanism by which lower NMI induces SG is in the geometry of representations. In particular, we find that NMI induces more parallelism in neural representations (i.e., input features coded in parallel neural vectors) of the model, a property related to the capacity of reasoning by analogy.
@@ 1 Introduction
A notable exception to this trend is the study of data complexity. Indeed, both in terms of patterns and scale, increasing training data complexity can improve SG in language models [38, 31, 51, 65, 1, 20]. Similarly, within the realm of vision, augmenting training data with translations has been shown to improve learned invariances [9]. Nevertheless, several questions remain open. First, although incrementing training data complexity does promote SG, which property (or properties) of data distribution subtends this promotion remains obscure. Second, despite the tremendous success and broad applicability of vision-language models [43, 41, 6, 42], whether data training properties can boost SG in these models has not been explored. Third, a systematic analysis of the effect of training data properties on internal representations and how these can promote SG is also missing. In other words, we lack a functional explanation of how data properties may promote SG.
In this work, we focus on the following research questions: ”Can SG be induced solely through changing properties of the training data? If so, how and why does this work?”. To answer these questions, we focus on a Multi-modal Masked Language Modeling (MMLM) setting and test for three different properties of the data: diversity, an increase in the cardinality of latent factors of the training distribution; burstiness, a limitation on the cardinality of latent factors in particular instances of the training data; and latent intervention, a targeted intervention on the value of a latent factor. We construct datasets that alter these properties by manipulating their generative latent factors and test how a model trained on these datasets generalizes to instances of unseen combinations of latent factors. Our experiments demonstrate that manipulating these properties the SG capability of a model.
Finally, to understand why these properties induce better SG, we perform a thorough set of experiments. We study the role of the Normalized Mutual Information (NMI) between latent attributes of a dataset and find that it correlates strongly with out-of-distribution (o.o.d) generalization. We then turn to the effect of data distributional properties on learned representations. Moreover, we analyze the representations’ geometry compared to baseline models. We find that the more diverse a dataset the more it induces disentanglement in the representations. Also, we analyze other geometrical properties of representations. There is currently a conjecture —based on empirical observations of biological neural representations— that the more parallel representations remain under a particular change in a latent attribute, the better they are at SG. Through our experiments, we find that NMI tends to produce representations that show more parallelism, suggesting a mechanism by which lower NMI datasets can induce SG. Thus, our contributions can be summarized as follows:
@@
  • We find that datasets with lower NMI between latent factors tend to produce models with better out-of-distribution performance. Moreover, we find that one mechanism by which they induce this behavior is by promoting the emergence of representations with more parallellism under changes in latent attributes.
@@ 2 Problem Formulation
We train a model on a similar task to CLEVR[32] to correctly respond to a textural query given contextual information (in our case, images) about a group of entities contained in it. In particular, queries that will give a partial description of the image (e.g., ”small red rubber sphere, large [?] metallic cylinder, etc.”) will always resolve to predict the value of one or more particular attributes (e.g., color in this case) of a particular entity (e.g., cube). See Figure 1 for an example input image, query, and expected answer. A model able to perform SG will achieve this over a test set with unseen combinations of these attributes during training. This is relevant because models tend to memorize combinations of attributes (which may prove spurious) during training rather than learn each one separately.
@@ 2.1 Task formalization
Altogether, we can define $\mathcal{D}=\{(c_{i},q_{i},y_{i}),i=1..N\}$ to be a dataset of $N$ samples, where $c_{i},q_{i},y_{i}$ refer to the context, queries and answers for the $i$-th element in the dataset. The task consists of predicting the correct $a_{i}$, given $c_{i}$ and $q_{i}$.
@@ 2.3 Distributional Properties of Training Data
We wish to study how three properties of $\mathcal{D}_{train}$ affect SG. Therefore, we look for distributional properties of $\mathcal{D}_{train}$ that might be able to break the link between $a$ and $b$. To achieve this, we propose three properties, give intuition on why they may work to achieve this goal, and define them.
@@ 2.3.1 Diversity
To disrupt the link between properties $a$ and $b$, we propose to increase the total number of values $a$ may take. Our hypothesis is that increasing the diversity of values of $a$ on the training set might make it harder for the model to memorize all combinations of $a$ and $b$. Thus, the model may prefer solutions that learn to detect $a$ and $b$ separately. Therefore, we propose diversity over attribute $a$ as the total cardinality of $a\in A$, that is $div(a)=|a_{i}|$. Over our experiments, we modulate diversity over $a$ in a dataset by altering the number of possible values $a$ may take in the training set.
@@ 3.1 Dataset
We create a dataset resembling the structure of the CLEVR dataset [32] by leveraging their generative code. Attributes for this dataset are $A=\{shape,color,material,size\}$. Our main goal is to predict the shape attribute systematically, while excluding certain combinations of shape-color from $\mathcal{D}_{train}$. There are three possible shapes: {sphere, cube, cylinder}. The context for this dataset is a 224 x 224 color image of 3 to 10 objects. Queries are defined as a textual description of the scene where the property of the object that is to be predicted is masked. Contextual information is also included in the query to make queries unambiguous. Our training set consists of 75.000 samples and the test set of 15.000.
@@ 3.4 Evaluation
We run each experiment for 3 seeds and report their mean and standard error. All metrics are related to performance on cubes and cylinders to correctly assess for generalization in objects with systematic differences to the training set.
Figure 1: (A) Example input for our task. Image shows objects with different properties, while query gives a textual description of the objects, while masking properties to be predicted. Sample query lists numbered objects from 1-3. (B) Diagram of the model during training. Context is provided as an image, while the query consists of text denoting information about the scene with some text related to properties of objects are masked. Both the context and query are entered into a Transformer Encoder, where it classifies for each masked token the most probable value of the masked property in the query.
@@ 4 Baseline Results on SG
Figure 2: Accuracy for predicting different properties for different data splits for the baseline model (8 colors). In-distribution performance for all tasks remains high, however out-of-distribution performance plummets for shape and color, suggesting the model is learning combinations of shape-color as features, instead of achieving systematic generalization. Unexpectedly, material and size also show a drop in out-of-distribution performance, even though the model has been exposed to all combinations of these attributes in training.
In-distribution performance remains similar to that of $\mathcal{D}_{train}$ for all tasks. However, delving into performance for $\mathcal{D}_{test-OOD}$ reveals three distinct generalization patterns. First, unexpectedly, the material and size show a decrease in $\mathcal{D}_{test-OOD}$ performance, even though the model has been exposed to all combinations of these attributes during training. Second, we observe a significant decrease in performance for the color task (44% drop) and shape (99% drop). This is especially bad for the shape task which achieves much worse than random performance (33% to predict between three shapes in $\mathcal{D}_{test-OOD}$). These results strongly suggest that the model faces challenges when performing SG. These results confirm previous findings that indicate DNNs encounter significant performance challenges when required to generalize systematically [37, 60, 34, 33]. Third, given that both shape and color are related systematically, a natural question arises: why is color less affected than shape? An explanation for this difference in performance between color and shape on $\mathcal{D}_{test-OOD}$ emerges from how image information is fed to the model. Due to the division of images into patches, the model can predict the color of a particular object solely based on the information of one patch. However, to predict the shape of an object, the model needs information from additional patches, making this ability harder to train and thus successfully achieve SG.
We turn to analyzing how modulating the amount of training data affects SG. A popular way of promoting generalization is to increase the amount of training data. Therefore, we train models on different percentages of $\mathcal{D}_{train}$. Results are shown in Figure 3(a). While in-distribution roughly stays the same, increasing the amount of training data does not result in better out-of-distribution generalization. Performance degrades as dataset size increases. This suggests that merely adding more in-distribution training data may not be a robust way of inducing a bias toward systematic generalization. Even worse, it seems to double down on whatever non-systematic bias the model is learning. We now alter the distribution of $\mathcal{D}_{train}$.
@@
Figure 3: Accuracy for the shape task for different amounts of training data for the baseline model (8 colors) vs model trained on 216 colors. (a) Increasing dataset size does not increase out-of-distribution performance but rather degrades it slightly. (b) With increased diversity much stronger out-of-distribution generalization is achieved, with models trained on only a quarter of the data severely outperforming the 8-color baseline.
5 Impact of Altering the Training Distribution on SG
We investigate the impact on SG of altering $\mathcal{D}_{train}$ by manipulating properties specifically tailored to perturbate the link between $a$ and $b$.
@@ 5.1 Diversity
Figure 4: Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for the color and shape tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the out-of-distribution split, while in-distribution only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
To increase diversity, we run experiments where we increase colors in $\mathcal{D}_{train}$ from 8 to 216. Figure 4 shows these results. We observe that augmenting color diversity improves SG on the shape task (see monotonic increase of green bars. This increase is of an absolute 89%, a staggering amount that reflects the difference between a model that does not generalize systematically to one that does. Unexpectedly, the performance on color prediction barely decreases in test-ID (around 6%) (Fig. 4(a)), in spite of the increased difficulty of the task given by a higher number of classes. Moreover, increased color diversity reduces the performance gap between $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ in the color task (Fig. 4(a)). Remarkably, out-of-distribution performance for the $material$ and $size$ tasks also increases significantly when augmenting colors. This suggests that increased diversity seems to be creating an inductive bias where at least color is being disassociated from shape, material and size. Moreover, this increase for material and size occurs rapidly with just 27 colors, while for shape this keeps increasing until reaching 216 colors.
Finally, we test whether increasing diversity affects the model’s sample efficiency. We proceed to train models on datasets with 216 colors with size $\{0.25,0.5,0.75\}$ of the original training size and compare them to the 8-color baseline at full training size. Results are shown in Figure 3(b). As can be seen, even training with just one quarter of the original training size with increased diversity produces much stronger out-of-distribution generalization than the baseline, while the gap between in-distribution and out-of-distribution remains relatively low.
@@ 5.3 Latent Intervention
Figure 6 shows results for modulating latent intervention over color. In our experiments, we can observe that increasing the degree of the perturbation we apply to the color of the images during training increases the model performance on the shape task by up to 15%, a non-trivial amount while retaining out-of-distribution performance on the color task. For the material and size tasks, latent intervention adds up to 2% in $\mathcal{D}_{test-OOD}$ performance, and the effect is reduced as the number of colors increases.
@@
Figure 6: Change in accuracy after applying latent intervention for the color and shape tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. X-Axis not at scale. Altering the color hue randomly during training allows the model to gain up to 15% more out-of-distribution accuracy over baseline.
@@ 5.4 Discussion
While all properties improve SG in the shape task, diversity clearly seems to have the greatest impact overall. However, we see that the other properties seem to be complementary to diversity. This suggests a positive answer to our research question “Can Systematic Generalization be induced solely through changing properties of the training data?”.
Surprisingly, the effects of altering these properties does not only affect performance in $\mathcal{D}_{test-OOD}$ for shape and color, but also material and size. This suggests that by default a model will tend to combine all latent features in the representation. We show that by especially targeting the spurious relation between latent factors through changes of the training distribution, this model behavior can be altered. We showed three ways in which this could be done with differing levels of success. This answers our second question: “How can we alter the training distribution to unduce SG?”. Finally, we move to answer why these properties affect SG. To do that, we will first study model capacity for SG.
@@ 6 Lack of Capacity does not explain SG
Given that diversity has shown to improve SG performance, we would like to gain a deeper understanding of why this is the case. Our first hypothesis is the following: adding more colors to the training distribution strains the capacity of the model such that it is forced to learn simpler learning solutions shape and color independently of each other. In other words, lack of capacity may force the model to find a more systematic solution. To test this, we test two cases: 1) We artificially restrict the capacity of a baseline Transformer model trained on 8 colors by limiting the size of the residual stream dimension ($d\in\{32,64,128\}$). 2) We augment the capacity of the model trained on 216 colors by increasing this dimension by $\{512,1024\}$.
@@
Figure 7: In-Distribution and Out-of-Distribution accuracy for models trained with different values for their hidden dimensions on a training set with 216 colors for the color and shape task. (a-b) When exposed to low diversity (8 colors) both metrics suffer as model capacity is reduced. (c-d) Both metrics increase as model capacity is increased. These results suggest that lack of capacity of the models is not the cause for achieving greater SG when augmenting the number of colors in the training set.

### v3

@@ Abstract
Deep neural networks (DNNs) struggle at systematic generalization (SG). Several studies have evaluated the possibility of promoting SG through the proposal of novel architectures, loss functions, or training methodologies. Few studies, however, have focused on the role of training data properties in promoting SG. In this work, we investigate the impact of certain data distributional properties, as inductive biases for the SG ability of a multi-modal language model. To this end, we study three different properties. First, data diversity, instantiated as an increase in the possible values a latent property in the training distribution may take. Second, burstiness, where we probabilistically restrict the number of possible values of latent factors on particular inputs during training. Third, latent intervention, where a particular latent factor is altered randomly during training. We find that all three factors significantly enhance SG, with diversity contributing an 89% absolute increase in accuracy in the most affected property. Through a series of experiments, we test various hypotheses to understand why these properties promote SG. Finally, we find that Normalized Mutual Information (NMI) between latent attributes in the training distribution is strongly predictive of out-of-distribution generalization. We find that a mechanism by which lower NMI induces SG is in the geometry of representations. In particular, we find that NMI induces more parallelism in neural representations (i.e., input features coded in parallel neural vectors) of the model, a property related to the capacity of reasoning by analogy. Our code is available at: https://github.com/fdelrio89/data-systematic
@@ 1 Introduction
A notable exception to this trend is the study of data complexity. Indeed, both in terms of patterns and scale, increasing training data complexity can improve SG in language models [44, 36, 62, 80, 1, 22]. Similarly, within the realm of vision, augmenting training data with translations has been shown to improve learned invariances [10]. Nevertheless, several questions remain open. First, although incrementing training data complexity does promote SG, the property (or properties) of data distribution that subtends this promotion remains obscure. Second, despite the tremendous success and broad applicability of vision-language models [49, 47, 7, 48], whether data training properties can boost SG in these models has not been explored. Third, a systematic analysis of the effect of training data properties on internal representations and how these can promote SG is also missing. In other words, we lack a functional explanation of how data properties may promote SG.
In this work, we focus on the following research questions: "Can SG be induced solely through changing properties of the training data? If so, how and why does this work?". To answer these questions, we focus on a Multi-modal Masked Language Modeling (MMLM) setting and test for three different properties of the data: diversity, an increase in the cardinality of latent factors of the training distribution; burstiness, a limitation on the cardinality of latent factors in particular instances of the training data; and latent intervention, a targeted intervention on the value of a latent factor. We construct datasets that alter these properties by manipulating their generative latent factors and test how a model trained on these datasets generalizes to instances of unseen combinations of latent factors. Our experiments demonstrate that manipulating these properties increases the SG capability of a model.
Finally, to understand why these properties enhance SG, we conduct thorough experiments. We study the role of the Normalized Mutual Information (NMI) between latent attributes of a dataset and find that it correlates strongly with out-of-distribution (OOD) generalization. We then turn to the effect of data distributional properties on learned representations. Moreover, we examine the representational geometry relative to baseline models. A prevailing conjecture, inspired by biological neural representations, suggests that the more parallel representations remain under a particular change in a latent attribute, the better they are at SG. Our experiments reveal that NMI influences this geometry. Thus, our contributions are summarized as follows:
  • We are the first to study the impact of training data properties and its effect on SG in the multimodal setting.
@@
  • While it is known that lower NMI between latent factors tend to produce models with better SG performance, we find that for datasets with the same level of NMI, performance differs greatly when altering dataset diversity. This suggests a different mechanism by which diversity increases OOD generalization.
  • We find that one mechanism by which lower NMI induces better SG is by promoting the emergence of representations with more parallellism under changes in latent attributes. To the best of our knowledge, this mechanism for SG has not been previously reported.
@@ 2 Problem Formulation
We train a model on a similar task to CLEVR [37] to correctly respond to a textural query given contextual information (in our case, images) about a group of entities contained in it. In particular, queries that will give a partial description of the image (e.g., "small red rubber sphere, large [?] metallic cylinder, etc.") will always resolve to predict the value of one or more particular attributes (e.g., color in this case) of a particular entity (e.g., cube). See Figure 1 for an example input image, query, and expected answer. A model able to perform SG will achieve this over a test set with unseen combinations of these attributes during training. This is relevant because models tend to memorize combinations of attributes (which may prove spurious) during training rather than learn each one separately.
@@ 2.1 Task formalization
Altogether, we define $\mathcal{D}=\{(c_{i},q_{i},y_{i}),i=1..N\}$ to be a dataset of $N$ samples, where $c_{i},q_{i},y_{i}$ refer to the context, queries and answers for the $i$-th element in the dataset. The task consists in predicting the correct $a_{i}$, given $c_{i}$ and $q_{i}$.
@@ 2.3 Distributional Properties of Training Data
We wish to study how three properties of $\mathcal{D}_{train}$ affect SG. Therefore, we look for distributional properties of $\mathcal{D}_{train}$ that might be able to break the link between $a$ and $b$. To achieve this, we propose three properties, define them, and give intuition on why they may work to achieve this goal.
@@ 2.3.1 Diversity
To disrupt the link between properties $a$ and $b$, we propose to increase the total number of values $a$ may take. We hypothesize that increasing the diversity of values of $a$ on the training set might make it harder for the model to memorize all combinations of $a$ and $b$. Thus, the model may prefer solutions that learn to detect $a$ and $b$ separately. Therefore, we propose diversity over attribute $a$ as the total cardinality of $a\in A$, that is $div(a)=|a_{i}|$. Over our experiments, we modulate diversity over $a$ in a dataset by altering the number of possible values $a$ may take in the training set.
@@ 3.1 Dataset
We create a dataset resembling the structure of the CLEVR dataset [37] by leveraging their generative code. Attributes for this dataset are $A=\{shape,color,material,size\}$. Our main goal is to predict the shape attribute systematically, while excluding certain combinations of shape-color from $\mathcal{D}_{train}$. There are three possible shapes: {sphere, cube, cylinder}. The context for this dataset is a 224 x 224 color image of 3 to 10 objects. Queries are defined as a textual description of the scene where the property of the object that is to be predicted is masked. Contextual information is also included in the query to make queries unambiguous. Our training set consists of 75,000 samples and the test set of 15,000.
@@ 3.4 Evaluation
We run each experiment for 3 seeds and report their mean and standard error. All metrics are related to performance on cubes and cylinders to correctly assess generalization in objects with systematic differences to the training set.
Figure 1: (A) Example input for our task. Image shows objects with different properties, while query gives a textual description of the objects while masking properties to be predicted. Sample query lists numbered objects from 1-3. (B) Diagram of the model during training. Context is an image, while the query is text describing the scene, with some text related to properties of objects being masked. Context and query enter into a Transformer that outputs its predictions for the values of the masked properties.
@@ 4 Baseline Results on SG
Figure 2: Accuracy of the baseline model (8 colors) for predicting different properties across data splits. While ID performance remains high, OOD performance drops sharply for shape and color, indicating the model learns shape-color combinations as features, rather than achieving SG. Surprisingly, material and size also show a drop in OOD performance, even though the model has been exposed to all combinations of these attributes in training.
ID performance remains similar to that of $\mathcal{D}_{train}$ for all tasks. However, delving into performance for $\mathcal{D}_{test-OOD}$ reveals three distinct generalization patterns. First, unexpectedly, the material and size show a decrease in $\mathcal{D}_{test-OOD}$ performance, even though the model has been exposed to all combinations of these attributes during training. Second, we observe a significant decrease in performance for the color task (44% drop) and shape (99% drop). This is especially bad for the shape task which achieves much worse than random performance (33% to predict between three shapes in $\mathcal{D}_{test-OOD}$). These results strongly suggest that the model faces challenges when performing SG. These results confirm previous findings that indicate DNNs encounter significant performance challenges when required to generalize systematically [43, 71, 39, 38]. Third, given that both shape and color are related systematically, a natural question arises: why is color less affected than shape? An explanation for this difference in performance between color and shape on $\mathcal{D}_{test-OOD}$ emerges from how image information is fed to the model. Due to the division of images into patches, the model can predict the color of a particular object solely based on the information of one patch. However, to predict the shape of an object, the model needs information from additional patches, making this ability harder to train and thus successfully achieve SG.
We turn to analyzing how modulating the amount of training data affects SG. A popular way of promoting generalization is to increase the amount of training data. Therefore, we train models on different percentages of $\mathcal{D}_{train}$. Results are shown in Figure 3(a). While ID roughly stays the same, increasing the amount of training data does not result in better OOD generalization. Performance degrades as dataset size increases. This suggests that merely adding more ID training data may not be a robust way of inducing a bias toward SG. Even worse, it seems to double down on whatever non-systematic bias the model is learning. We now alter the distribution of $\mathcal{D}_{train}$.
@@
Figure 3: Accuracy on the shape task with varying training data sizes for models trained on 8 colors (baseline) vs. 216 colors. (a) Increasing dataset size does not improve OOD performance and slightly degrades it. (b) Greater diversity significantly enhances OOD generalization, with models trained on just a quarter of the data severely outperforming the 8-color baseline.
5 Altering the Training Distribution for SG
We investigate the impact on SG of altering $\mathcal{D}_{train}$ by modifying properties that disrupt the link between $a$ and $b$.
@@ 5.1 Diversity
Figure 4: Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for the color and shape tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the OOD split, while ID only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
To increase diversity, we run experiments in which we increase the colors in $\mathcal{D}_{train}$ from 8 to 216. As shown in Figure 4, greater color diversity improves SG on the shape task (see the monotonic increase of green bars). An absolute improvement of $89\%$, a staggering amount that reflects the difference between a model that does not generalize systematically and one that does. Surprisingly, the test-ID color prediction drops only slightly ($\sim 6\%$) (Fig. 4(a)), despite the increased difficulty of an increased class count. Moreover, higher color diversity reduces the performance gap between $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ in the color task (Fig. 4(a)). Remarkably, OOD performance for the $material$ and $size$ tasks also increases significantly when adding colors. This suggests that the increase in diversity appears to be creating an inductive bias where at least color is being disassociated from shape, material, and size. Moreover, this increase for material and size occurs rapidly with just 27 colors, while for shape, this keeps increasing until reaching 216 colors.
Finally, we assess whether increased diversity affects sample efficiency. We train models on datasets with 216 colors of $\{0.25,0.5,0.75\}$ the original training size and compare them to the 8-color baseline at full size. The results are shown in Figure 3(b). Remarkably, even with just $25\%$ of the data with increased diversity, produces much stronger OOD generalization than the baseline, while the gap between ID and OOD remains relatively low.
@@ 5.3 Latent Intervention
Figure 6 shows results for modulating latent intervention over color. In our experiments, we can observe that increasing the degree of the perturbation we apply to the color of the images during training increases the model performance on the shape task by up to 15%, a non-trivial amount while retaining OOD performance on the color task. For the material and size tasks, latent intervention adds up to 2% in $\mathcal{D}_{test-OOD}$ performance, and the effect is reduced as the number of colors increases.
@@
Figure 6: Change in accuracy after applying latent intervention for the color and shape tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. X-Axis not at scale. Altering the color hue randomly during training allows the model to gain up to 15% more OOD accuracy over baseline.
@@ 5.4 Discussion
While all properties improve SG in the shape task, diversity clearly seems to have the greatest impact overall. However, we see that the other properties seem to be complementary to diversity. This suggests a positive answer to our research question, “Can SG be induced solely through changing properties of the training data?".
Surprisingly, altering these properties affects not only performance on $\mathcal{D}_{test-OOD}$ for shape and color, but also material and size. This suggests that, by default, a model tends to entangle all latent features in its representation. We demonstrate that modifying the training distribution to specifically target spurious relations between latent factors can alter this behavior. We explored three approaches with varying success, addressing our second question: “How can we modify the training distribution to induce SG?". Finally, we turn to understanding why these properties influence SG, beginning with an analysis of model capacity.
@@ 6 Lack of Capacity does not explain SG
Given that diversity has shown to improve SG performance, we would like to gain a deeper understanding of why this is the case. Our first hypothesis is the following: adding more colors to the training distribution pressures the capacity of the model, forcing it to learn simpler solutions where shape and color are represented independently. In other words, lack of capacity may force the model to find a more systematic solution. To test this, we test two cases: (1) We artificially restrict the capacity of a baseline Transformer model trained on 8 colors by limiting the size of the residual stream dimension ($d\in\{32,64,128\}$). (2) We augment the capacity of the model trained on 216 colors by increasing this dimension by $\{512,1024\}$.
@@
Figure 7: ID and OOD accuracy for models trained with varying hidden dimensions on a 8 (a-b) and 216 (c-d) color dataset for the color and shape task. (a-b) With low diversity (8 colors) both metrics drop as capacity decreases. (c-d) With higher diversity, both slightly improve as capacity increases. These suggest that a capacity-bottleneck is not the cause for achieving greater SG when augmenting the number of colors in the training set.